OpenStudy (anonymous):

Write and then solve the differential equation for the statement: "The rate of change of y with respect to x is inversely proportional to the square root of y."

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\sqrt{y}\]

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

Have I written out the prompt properly?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \frac{dy }{dx} =k\times\sqrt{y} }\) that would be true, if you say "varies"

OpenStudy (solomonzelman):

if you say "inversely varies" then \(\large\color{#000000 }{ \displaystyle \frac{dy }{dx} =k\times(1/\sqrt{y}) }\)

OpenStudy (solomonzelman):

(I think that is how it is....)

OpenStudy (anonymous):

Hmm I don't "varies" in the question so would I still include a constant, k?

OpenStudy (solomonzelman):

proportional=varies

OpenStudy (solomonzelman):

inversely proportional = varies inversely

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \frac{dy }{dx} =k\times(1/\sqrt{y}) }\) \(\large\color{#000000 }{ \displaystyle \sqrt{y}\frac{dy }{dx} =k }\) \(\large\color{#000000 }{ \displaystyle \color{#ff0000}{\int}\sqrt{y}\frac{dy }{dx} \color{#ff0000}{~dx}=\color{#ff0000}{\int}k\color{#ff0000}{~dx} }\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \int\sqrt{y}~dy=\int k~dx }\)

OpenStudy (solomonzelman):

go from there ...

OpenStudy (solomonzelman):

(Note, you need to solve this equation for y, and don't forget about the integration constant C)

OpenStudy (solomonzelman):

I got to go...

OpenStudy (anonymous):

I'll try to figure it out. Thanks

OpenStudy (anonymous):

I got \[y=(\frac{ 3 }{ 2}kx)^{2/3}+C\]

OpenStudy (anonymous):

@ganeshie8 @freckles Can any of you confirm for me?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@baru

OpenStudy (baru):

i think you have gone wrong in the integration part. can you tell the rule for integrating \(x^n\)?

OpenStudy (anonymous):

The main thing that has me confused is C, I am not sure if they cancel each other out when you have something like this: \[x+C=12x+C\]

OpenStudy (anonymous):

@baru integration is correct, checked using wolfram

OpenStudy (baru):

no..they do not cancel

OpenStudy (anonymous):

So I would still have \[x=12x+C\] When moving C to the other side?

OpenStudy (baru):

yes, well actually, if it confuses, you.. this is how it is actually supposed to be written \[x+C_1=12x+C_2\]

OpenStudy (baru):

\[x=12x+(C_2-C_1)\]

OpenStudy (baru):

since c2 and C1 are both arbitrary constants, we replace C2-C1 with just C

OpenStudy (baru):

i seem to be getting\[\frac{2}{3}y^{2/3}=kx +c\\y^{2/3}=\frac{3}{2}kx +C\\y=(\frac{3}{2}kx +C)^{3/2}\]

OpenStudy (anonymous):

should be y=(3/2kx + C)^2/3

OpenStudy (whpalmer4):

@baru \[\int y^n\ dy = \frac{1}{n+1}y^{n+1} + C \]\[\int \sqrt{y}\ dy = \int y^{1/2}\ dy = \frac{1}{1+\frac{1}{2}} y^{1+\frac{1}{2}} + C = \frac{2}{3}y^{3/2} +C\] your exponent is incorrect, should be \[\frac{2}{3}y^{3/2} = kx + c\]\[y^{3/2} = \frac{3}{2}kx + C\]\[y = (\frac{3}{2}kx + C)^{2/3}\]

OpenStudy (baru):

yes...my mistake :)